# Risk-Based Capacity Accreditation of Resource- Colocated Large Loads in Capacity Markets

### v1

@@ Abstract
We study capacity accreditation of resource-colocated large loads, defined as large demands such as data center and manufacturing loads colocated with behind-the-meter generation and storage resources, synchronously connected to the bulk power system, and capable of participating in the wholesale electricity market as an integrated unit. Because the qualified capacity of a resource portfolio is not equal to the sum of its individual resources’ qualified capacities, we propose a novel risk-based capacity accreditation framework that evaluates the collective contribution to system reliability. Grounded in the effective load carrying capability (ELCC) metric, the proposed capacity accreditation employs a convex optimization engine that jointly dispatches colocated resources to minimize reliability risk. We apply the developed methodology to a hydrogen manufacturing facility with colocated renewable generation, storage, and fuel cell resources.
@@ I Introduction
The global electric power sector is experiencing an unprecedented surge in electricity demand. With an anticipated 3.5% compound annual growth rate over the next 25 years, electricity consumption in 2050 is projected to exceed 2.5 times the 2023 level of 25,000 TWh. The top three sectors that fuel the demand growth are the “large loads” from data centers, hydrogen production, and electrified transportation [McKinsey:24].
@@
A promising but underdeveloped approach to support resource adequacy amid rising large load is Resource Colocation [^1] Both colocation and collocation appear in the literature to describe multiple resources located together physically, with colocation emphasizing the strategic placement and coordination to achieve operational or economic objectives. Some publications make a distinction between colocated and hybrid resources depending on the metering and the coordination among colocated resources. This paper does not make such distinction.. In this context, a resource-colocated large load is a large load paired with behind-the-meter generation and storage resources that are synchronized with the bulk power system and participate in the wholesale market as a single entity. Fig. 1 illustrates a renewable-colocated hydrogen manufacturing facility with grid-connected renewable generation, an electrolyzer to produce hydrogen sold in a hydrogen market, and a hydrogen storage tank with the ability to generate power via fuel cell from stored hydrogen. Such a colocated hydrogen production resource can participate in the energy, ancillary service, and capacity markets.
@@ I-A Related Work
Capacity accreditation is a cornerstone of deregulated capacity market design. It maps a resource’s nameplate capacity to its qualified capacity by accounting for resource-level and system-level reliability constraints, such as loss of load probability (LOLP), loss of load expectation (LOLE), and expected unserved energy (EUE).
Traditional capacity accreditation—most notably through the computation of ELCC [Garver:1966]—quantifies a resource’s capacity value as the additional system load that can be supported while maintaining a target reliability level. Over time, a rich set of ELCC approximations and extensions has been developed for diverse resource types [Milligan&etal:16EJ], including colocated or hybrid resources [Schlag&etal:20E3Rpt, Stenclik&Goggin&Ela&Ahlstrom:ESIG22Rpt, Ericson&etal:22NREL, Wen&Song:2023]. System operators have also implemented rule-based heuristics [PJM_Manual20A, zhao2024] that use predetermined priorities to simplify ELCC calculations. Most existing approaches, however, evaluate specific operational strategies of utilizing system and colocated resources. Such approaches may either overestimate or underestimate a resource’s actual contribution to adequacy, whereas the risk-based methodology presented here emphasizes the optimal utilization of colocated resources.
There is a substantial literature on the grid integration of large loads; see a recent review in [Frick&Srinivasan:25LBNL]. The literature on colocated loads is limited but emerging. Particularly relevant here are studies of colocated hydrogen production and data centers in energy and ancillary service markets [Glenk&Reichelstein:19Nature, Wu&etal:Energy22, Li&etal:25arxiv, Li&etal:26PESIM]. This paper is perhaps the first to examine the capacity accreditation of resource-colocated large loads in a capacity market.
@@ I-B Summary of Contribution
The main technical contribution of this paper is threefold. First, we develop a risk-minimizing approach to evaluate the ELCC of resource-colocated large loads. Since ELCC fundamentally depends on the underlying reliability assessment, we propose in Sec. III a convex optimization framework that evaluates system reliability through a reliability loss minimization formulation, in contrast to the rule-based heuristics commonly used in the literature and in practice. This formulation renders existing heuristic practices as feasible (but generally suboptimal) solutions to the proposed optimization. Second, in Sec. IV, we introduce a scenario-based procedure for ELCC evaluation that accounts for stochastic resource availability and load variability. Finally, in Sec. V, we present a novel application of the proposed capacity accreditation methodology to a hydrogen manufacturing system colocated with on-site renewable generation, long-duration hydrogen storage, and fuel cell-based power production.
@@ II-B ELCC for Colocated Resources
By its standard definition [Garver:1966], ELCC quantifies the additional load [^2] Unlike flexible demand, which represents controllable demand-side resources, load refers to the system’s inflexible electricity demand. that a power system can supply while maintaining the same target reliability after adding a new resource. The calculation proceeds by first evaluating the reliability of a baseline system, then adding the resource and increasing the load until the reliability returns to the baseline level. The resulting load increase is the ELCC of the resource.
@@ III-A Reliability Loss Minimization
$\bm{P}_{t}\coloneqq\left\{\begin{aligned} &\{P^{\mbox{\tiny\sf CH}}_{i,t},\,P^{\mbox{\tiny\sf DIS}}_{i,t},\,e_{i,t}\}_{i\in[n_{\text{\sf STO}}]},\\ &\{P^{\mbox{\tiny\sf FLEX}}_{i,t},\,P^{\mbox{\tiny\sf RED}}_{i,t}\}_{i\in[n_{\text{\sf FLEX}}]},\\ &\{\bm{P}^{\mbox{\tiny\sf COL}}_{i,t}\}_{i\in[n_{\text{\sf COL}}]},P_{t}^{\mbox{\tiny\sf CUR}},\,P_{t}^{\mbox{\tiny\sf UE}}\end{aligned}\right\}.$
@@
  • For LOLE minimization:
@@
We define $G_{t}$ as the total power supply at time $t$, including generation, storage discharge, and colocated output $g(\bm{P}_{t}^{\mbox{\tiny\sf COL}})$. $L_{t}$ denotes the total demand, comprising inflexible load and the actual consumption of flexible-demand and colocated resources $l(\bm{P}_{t}^{\mbox{\tiny\sf COL}})$. For notational convenience, $g(\cdot)$ and $l(\cdot)$ denote the aggregate colocated output and consumption. Equations (1b)–(1d) enforce power balance, with shortfalls represented by unserved energy $P_{t}^{\mbox{\tiny\sf UE}}$ and surpluses by curtailment $P_{t}^{\mbox{\tiny\sf CUR}}$. Equations (1e)–(1h) specify storage operations, including charge/discharge limits and state-of-charge dynamics. Flexible-demand constraints (1i)–(1k) ensures that demand reductions occur only under supply shortages and apply solely to covering the shortfall, consistent with industry practices [PJM_Manual20A]. (1l) defines a generic constraint set representing the operational characteristics of colocated resources.
@@ IV Risk-based Capacity Accreditation
Following Garver’s classical definition [Garver:1966], system load is represented by the annual peak load $Q^{\mbox{\tiny\sf D}}$. An increase in load by $\Delta$ corresponds to raising the peak load by $\Delta$ while scaling the load profile proportionally:
@@
In capacity accreditation applications, system operators often compute the ELCC of a resource relative to different benchmarks or reference units. Examples include defining ELCC as the equivalent increase in load that can be supplied at the target reliability level, as the capacity of a perfectly reliable generator, or as that of a reference unit with a specified forced outage rate [NERC_2011]. The proposed approach is general and can be applied to any of these benchmark-based calculations.
@@ V ELCC of Resource-Colocated Hydrogen
We studied a hydrogen manufacturing system colocated with renewable generation, flexible-demand electrolyzers, and hydrogen storage, as illustrated in Fig. 1. The electrolyzer can modulate its power consumption within a specified operating range and convert electricity into hydrogen, which can either be sold directly or stored in a hydrogen tank. The ability to adjust power drawn from the grid makes the hydrogen producer a flexible-demand resource.
@@
We applied the proposed approach to evaluate the ELCC of the colocated resource, where the colocated wind capacity was 10 MW, the nominal electrolyzer demand was 8 MW with up to 50% demand response capability, and the stored hydrogen could generate up to 5 MWh of electricity. The analysis was performed on a calibrated baseline power system with a 650 MW peak load, 640 MW of conventional generation capacity, 300 MW of wind, 100 MW of solar, and a total of 700 MWh of energy storage. The load and renewable generation profiles were derived from publicly available NYISO real-time data [NYISO:Data].
Table I presents the ELCC of the colocated hydrogen system as its installed capacity scales, for an annual EUE target of 100 MWh. We compared a heuristic accreditation approach, which follows system operator scheduling rules to determine system reliability [PJM_Manual20A], with the proposed optimization-based method. Across all capacity scaling factors, the proposed method accredited more capacity than the heuristic approach, indicating that the optimization-based reliability model captured the adequacy contribution of the colocated resource more effectively.
TABLE I: ELCC of the colocated resource under heuristic and proposed methods as the capacity ratio varies.
Capacity scaling factor 1.0 1.2 1.4 1.6 1.8 2.0
Heuristic (MW) 5.28 6.34 7.39 8.44 9.49 10.54
Proposed method (MW) 5.35 6.41 7.47 8.52 9.57 10.60
Figure 3 shows the ELCC values for different installed capacities under three reliability targets. The data labels indicate the ELCC as a percentage of the installed capacity for the EUE = 200 MWh case. We observed that this percentage declined modestly as capacity increased, reflecting diminishing marginal ELCC in this setting.
Figure 3: ELCC versus installed capacity for the colocated resource under different EUE targets.
@@ VI Conclusions
Colocating large loads with on-site renewable and storage resources is a promising solution for meeting demand surges from large loads. Resource colocation not only reduces operating costs and enhances reliability for the large loads themselves, but also benefits power system operation by reducing peak demand. This paper demonstrates the under-explored benefits of resource colocation to resource adequacy through capacity markets. Although it is known that the ELCC of a resource portfolio is not additive, how to compute ELCC accurately for colocated resources has not been formalized to date. This paper fills this gap with a risk-minimization approach that generalizes existing solutions and applies to colocated data center and manufacturing large loads.

### v2

@@ Abstract
We study capacity accreditation of resource-colocated large loads, defined as large demands such as data center and manufacturing loads colocated with behind-the-meter generation and storage resources, synchronously connected to the bulk power system, and capable of participating in the wholesale electricity market as an integrated unit. Because the accredited capacity of a resource portfolio is not equal to the sum of its individual resources’ capacity values, we adopt a risk-based capacity accreditation framework to evaluate the combined reliability contribution of colocated resources. Grounded in the effective load carrying capability (ELCC) metric, the proposed capacity accreditation employs a convex optimization engine that jointly dispatches colocated resources to minimize reliability risk. We apply the developed methodology to a hydrogen manufacturing facility with colocated renewable generation, storage, and fuel cell resources.
@@ I Introduction
The global electric power sector is experiencing an unprecedented surge in electricity demand. With an anticipated 3.5% compound annual growth rate over the next 25 years, electricity consumption in 2050 is projected to exceed 2.5 times the 2023 level of 25,000 TWh. The top three sectors that fuel the demand growth are the “large loads” from data centers, hydrogen production, and electrified transportation [8].
@@
A promising but underdeveloped approach to support resource adequacy amid rising large load is Resource Colocation. [^1] Both colocation and collocation are used in the literature to describe physically colocated resources, with colocation often emphasizing coordinated operation for operational or economic objectives. Some studies distinguish between colocated and hybrid resources based on metering and coordination; this paper does not make such a distinction. In this context, a resource-colocated large load is a large load paired with behind-the-meter generation and storage resources that are synchronized with the bulk power system and participate in the wholesale market as a single entity. Fig. 1 illustrates a renewable-colocated hydrogen manufacturing facility with grid-connected renewable generation, an electrolyzer to produce hydrogen sold in a hydrogen market, and a hydrogen storage tank with the ability to generate power via fuel cell from stored hydrogen. Such a colocated hydrogen production resource can participate in the energy, ancillary service, and capacity markets.
@@ I-A Related Work
Capacity accreditation is typically formulated within a risk-based framework. This framework maps a resource’s nameplate capacity to its accredited value by accounting for system reliability risk. Accredited capacity is most commonly quantified using metrics such as ELCC or marginal reliability impact (MRI) [4, 17].
Most existing studies focus on the accreditation of individual resources [13, 1, 3]. With the ongoing evolution of power systems, increasing attention has been paid to portfolio effects in capacity accreditation. Several studies have examined the capacity value of colocated and hybrid resources, recognizing interactions among different resource types [12, 14, 2, 15]. In these works, the combined capacity value of multiple resources is assessed; however, the resources are generally not modeled as a single entity with internally coordinated dispatch decisions, nor are the internal operational constraints of colocated resources explicitly represented. While the literature on colocated loads remains limited, related studies on colocated hydrogen production and data centers in energy and ancillary service markets have begun to emerge [5, 16, 7, 6]. To the best of our knowledge, this paper is perhaps the first to examine the capacity accreditation of resource-colocated large loads in a capacity market.
Methodologies for capacity accreditation of colocated resources remain insufficiently developed. A commonly proposed portfolio-level approach is to sum the accredited capacities of individual components [14, 2], which can yield inaccurate results due to interactions among resources. An alternative approach applies the risk-based accreditation framework at the aggregate level by dispatching the colocated resource according to extensions of existing scheduling rules [14, 2, 15]. In the literature, such scheduling is typically driven by cost minimization or profit maximization in electricity markets [13]. To accurately capture reliability contributions, some studies adopt dispatch strategies that directly minimize reliability risk [1, 3], which are most closely aligned with our approach. However, these methods often incur substantial computational burdens and therefore rely on simplifying assumptions—such as homogeneous storage efficiencies—to maintain tractability. Rule-based scheduling heuristics, such as those adopted in system operator practices [11], are computationally efficient, but as approximations of optimal coordination, they do not guarantee accurate capacity accreditation.
@@ I-B Summary of Contribution
The main technical contributions of this paper are threefold. First, we propose a convex optimization-based computation engine that assesses system reliability through a reliability loss minimization formulation, in contrast to the rule-based heuristics commonly used in the literature and in practice. This formulation interprets existing heuristic practices as feasible, though generally suboptimal, solutions to the proposed optimization. Second, in Sec. IV, we introduce a scenario-based procedure for ELCC evaluation that accounts for stochastic resource availability and load variability. Finally, in Sec. V, we present a novel application of the proposed capacity accreditation methodology to a hydrogen manufacturing system colocated with on-site renewable generation, long-duration hydrogen storage, and fuel cell-based power production.
@@ II-B ELCC for Colocated Resources
By its standard definition [4], ELCC quantifies the additional load [^2] Unlike flexible demand, which represents controllable demand-side resources, load refers to the system’s inflexible electricity demand. that a power system can supply while maintaining the same target reliability after adding a new resource. The calculation proceeds by first evaluating the reliability of a baseline system, then adding the resource and increasing the load until the reliability returns to the baseline level. The resulting load increase is the ELCC of the resource.
@@ III-A Reliability Loss Minimization
$\bm{P}_{t}\coloneqq\left\{\begin{aligned} &\{P^{\mbox{\tiny\sf CH}}_{i,t},\,P^{\mbox{\tiny\sf DIS}}_{i,t},\,e_{i,t}\}_{i\in[n_{\text{\sf STO}}]},\\ &\{P^{\mbox{\tiny\sf FLEX}}_{i,t},\,P^{\mbox{\tiny\sf RED}}_{i,t}\}_{i\in[n_{\text{\sf FLEX}}]},\\ &\{\bm{P}^{\mbox{\tiny\sf COL}}_{i,t}\}_{i\in[n_{\text{\sf COL}}]},P_{t}^{\mbox{\tiny\sf CUR}},\,P_{t}^{\mbox{\tiny\sf UE}}\end{aligned}\right\},$
@@
  • For loss of load expectation (LOLE) minimization:
@@
We define $G_{t}$ as the total power supply at time $t$, including generation, storage discharge, and colocated output $g(\bm{P}_{t}^{\mbox{\tiny\sf COL}})$. $L_{t}$ denotes the total demand, comprising inflexible load and the actual consumption of flexible-demand and colocated resources $l(\bm{P}_{t}^{\mbox{\tiny\sf COL}})$. For notational convenience, $g(\cdot)$ and $l(\cdot)$ denote the aggregate colocated output and consumption. Equations (1b)–(1d) enforce power balance, with shortfalls represented by unserved energy $P_{t}^{\mbox{\tiny\sf UE}}$ and surpluses by curtailment $P_{t}^{\mbox{\tiny\sf CUR}}$. Equations (1e)–(1h) specify storage operations, including charge/discharge limits and state-of-charge dynamics. Flexible-demand constraints (1i)–(1k) ensures that demand reductions occur only under supply shortages and apply solely to covering the shortfall, consistent with industry practices [11]. (1l) defines a generic constraint set representing the operational characteristics of colocated resources.
@@ IV Risk-based Capacity Accreditation
Following Garver’s classical definition [4], system load is represented by the annual peak load $Q^{\mbox{\tiny\sf D}}$. An increase in load by $\Delta$ corresponds to raising the peak load by $\Delta$ while scaling the load profile proportionally:
@@
In capacity accreditation applications, system operators often compute the ELCC of a resource relative to different benchmarks or reference units. Examples include defining ELCC as the equivalent increase in load that can be supplied at the target reliability level, as the capacity of a perfectly reliable generator, or as that of a reference unit with a specified forced outage rate [10]. The proposed approach is general and can be applied to any of these benchmark-based calculations.
@@ V ELCC of Resource-Colocated Hydrogen
System description: We studied a hydrogen manufacturing system colocated with renewable generation, electrolyzers, and hydrogen storage with fuel cell, as illustrated in Fig. 1. The electrolyzer can modulate its power consumption within a specified operating range and convert electricity into hydrogen, which can either be sold directly or stored in a hydrogen tank. The ability to adjust power drawn from the grid makes the hydrogen producer a flexible-demand resource.
@@
Case study setup: In the case study, the colocated wind capacity was 14 MW, the nominal electrolyzer demand was 14 MW with up to 30% demand response capability, and the stored hydrogen could generate up to 11.2 MW of electricity. The analysis was performed on a calibrated baseline power system with a 650 MW peak load, 640 MW of conventional generation capacity, 300 MW of wind, 100 MW of solar, and a total of 700 MWh of energy storage. The load and renewable generation profiles were derived from publicly available NYISO real-time data [9].
ELCC evaluation under different reliability metrics: We evaluated resource-specific and system-level ELCC for the colocated hydrogen system under LOLE- and EUE-based reliability criteria. Here, the qualified capacity was defined as the nameplate capacity for renewable, the reducible MW of flexible demand, the power rating for storage, and the sum of these for the colocated system. All reported percentages are relative to these definitions.
As shown in Fig. 3 (left), under LOLE minimization, wind achieved an ELCC of 7.8 kW, representing less than 0.1% of its qualified capacity when participating as a single resource in the capacity market. Flexible demand and storage reached ELCCs of 4.2 MW (100%) and 9.7 MW (86.7%), respectively. The colocated system attained 16.4 MW (55.9%), exceeding the sum of its individual components.
Figure 3: ELCC of resources evaluated under LOLE- and EUE-based reliability criteria.
Fig. 3 (right) presents ELCC values based on EUE. The renewable achieved an ELCC of 1.4 MW (10.1%), while demand response and storage were credited 4.2 MW (100%) and 9.0 MW (80.2%), respectively. The colocated hydrogen system received 15.6 MW (53.2%), again surpassing the sum of its individual components. These results demonstrate that the ELCC of a colocated system cannot be obtained by simply summing the ELCCs of its constituent resources.
Figure 4: ELCC differences of the colocated resource between the heuristic and proposed methods under different baseline system peak load levels.
ELCC differences between heuristic and proposed methods: Fig. 4 compares the ELCC of the colocated hydrogen system under a fixed scheduling priority defined by the operator [11] with the ELCC obtained from the proposed risk-minimization procedure. In the simulation, the baseline system load profile was uniformly scaled to achieve different peak load levels, while the installed generation capacity remained unchanged. The bars show the relative error, $\eta$, of the heuristic ELCC evaluation based on the EUE reliability criterion. Specifically, given the heuristic ELCC estimate ($\hat{\mbox{\small ELCC}}$) and the ELCC by definition ($\mbox{{\small ELCC}}^{*}$) obtained via risk minimization,
$\eta=(\hat{\mbox{\small ELCC}}-\mbox{{\small ELCC}}^{*})/(\mbox{{\small ELCC}}^{*}).$
We observed that, except at a system peak load of 500 MW, the heuristic method consistently underestimated ELCC, with errors ranging from just over 2% to more than 10%. Such discrepancies could result in substantial monetary losses for the colocated resource.
Fig. 4 also shows the EUE computed using the heuristic and the proposed approaches. Across the full range of peak loads, the heuristic method consistently overestimated EUE. Both methods exhibited a convexly increasing trend, reflecting the growing difficulty of meeting demand during scarcity hours under more stressed system conditions.
@@ VI Conclusions
Colocating large loads with on-site renewable and storage resources is a promising solution for managing grid demand surges. This paper highlights the underexplored benefits of colocation for resource adequacy in capacity markets. We propose an efficient risk-minimization approach for accurately evaluating the ELCC of colocated resources, generalizing existing methods and applicable to large loads such as data centers and manufacturing facilities. Our case study demonstrates that simple approaches—such as summing individual ELCCs or using heuristic estimates—can produce significant errors when applied to colocated systems, underscoring the need for rigorous evaluation methods.
Acknowledgments
The authors thank Prof. Le Xie and Dr. Tongxin Zheng for insightful discussions on capacity accreditation methodologies.